OpenStudy (anonymous):

What about this one? :/ Find the following infinite sum, if it exists. (LOOK AT RESPONSE) Thank you!

OpenStudy (anonymous):

\[\sum_{n-1}^{\infty}2\left(\begin{matrix}-1 \\ 2\end{matrix}\right)^{n-1}\]